X Company uses the high-low method to predict monthly overhead costs. The following were May and September cost and activity results:

OH Cost Production
May $9,398   2,700     
September $14,897   5,050     

If December production is expected to be 3,900 units, what are expected total fixed overhead costs in December [round unit costs to two decimal places]?
Answer #1

Answer:

Variable cost per unit = ($14,897 - $9,398) / (5,050 - 2,700) = $2.34

Fixed cost = $14,897 - (5,050 × $2.34) = $3,080

Or

Fixed cost = $9,398 - (2,700 × $2.34) = $3,080

Since fixed cost for all level of production is same, for December:-

Expected total fixed overhead costs in December = $3,080